OpenStudy (anonymous):

What are the points of discontinuity? y=((x+1)(x-5))/(x^2+4x+3)

OpenStudy (lυἶცἶ0210):

Have you tried factoring?

OpenStudy (anonymous):

Like this? y=(x-5)/(x+3)

OpenStudy (lυἶცἶ0210):

Yes, and what you elimated is where the discontinuity is

OpenStudy (lυἶცἶ0210):

Just plug that x coordinate in to find y :)

OpenStudy (anonymous):

Oh! Okay that makes a lot more sense. So is it x=5 and x=-3?

OpenStudy (lυἶცἶ0210):

Actually, when you factored, you got a common factor of x+1 right? Well solve that, x+1=0. So x= -1.

OpenStudy (freckles):

there is continuity when either factor of x^2+4x+3 is zero

OpenStudy (freckles):

you basically need to solve x^2+4x+3=0

OpenStudy (freckles):

you are already found one number such that y=f(x) is discontinuous

OpenStudy (freckles):

you just need to find the other

OpenStudy (freckles):

the types of discontinuity are different the x=-1 you have a hole because (x+1)/(x+1)=1 when x isn't -1 the other discontinuity you have a vertical asymptote

OpenStudy (lυἶცἶ0210):

Yea, whoops, sorry about the mix up ^

OpenStudy (freckles):

@FutureAnesthetist did you find the other discontinity?

OpenStudy (anonymous):

Yes, I got it! Thank you guys!

OpenStudy (anonymous):

Let me double-check real quick, it's x=-3, and x-1?

OpenStudy (freckles):

sounds beautiful